# Exceptional point and hysteresis in perturbations of Kerr black holes

### v2

@@ Abstract
We employ the isomonodromic method to study linear scalar massive perturbations of Kerr black holes for generic scalar masses $M\mu$ and generic black hole spins $a/M$. A degeneracy between the longest-living quasinormal mode and the first overtone is found for $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$. We show that the longest-living mode and the first overtone change continuously into each other as we vary the parameters around the point of degeneracy, providing evidence for the existence of a geometric phase around an exceptional point. We interpret our findings through a thermodynamic analogy.
@@
$\partial_{u}\left[(1-u^{2})\partial_{u}S\right]+\left(\frac{m^{2}}{1-u^{2}}-a^{2}c^{2}u^{2}-\Lambda\right)S=0,$ (3)
@@
$\begin{gathered}\theta_{0}=-m,\quad\theta_{t}=m,\quad\theta_{\star}=0,\quad\alpha=2\sqrt{\mu^{2}-\omega^{2}},\\ t=-2\alpha,\quad tc_{t}=\lambda+(1-m)\alpha+(\omega^{2}-\mu^{2})a^{2},\end{gathered}$ (5)
@@
$\begin{gathered}\theta_{0}=\frac{-i}{2\pi T_{-}}\left(\omega-m\Omega_{-}\right),\quad\theta_{t}=\frac{i}{2\pi T_{+}}\left(\omega-m\Omega_{+}\right),\\ \theta_{\star}=\frac{4M(\mu^{2}-2\omega^{2})}{\alpha},\quad t=2\alpha(r_{+}-r_{-}),\\ tc_{t}=\lambda+r_{+}^{2}\mu^{2}-(3a^{2}+r_{-}^{2}+3r_{+}^{2})\omega^{2}\\ \hskip 31.2982pt+\frac{\alpha}{2}\left[r_{-}-r_{+}-2iam+2i(a^{2}+r_{+}^{2})\omega\right]\\ \hskip 24.18501pt+\frac{2M(\mu^{2}-2\omega^{2})}{\alpha}-\frac{2M(\mu^{2}-2\omega^{2})^{2}}{\alpha^{2}}.\end{gathered}$ (6)
@@
The parameter $\sigma$ determined this way solves (7b). In fact, it parametrizes the complex phase of (8) as one continues it analytically over a path encompassing the singular points $z=0$ and $z=t$. In contrast, the interpretation of the monodromy parameter $\eta$ is quite involved [85]. For our purposes, we can just define the parameter by formally inverting the zero condition for $\tau_{V}$ (7a):
@@
where we have used $\theta_{0}$ and $\theta_{t}$ as defined in (5), and related the parameter $\sigma$ to the angular quantum number $\ell$. This equation relates the eigenvalues $\lambda$ and $\omega$ associated with the QNM problem. For $M\alpha\ll 1$, the expansion of $\lambda=\lambda(\omega)$ coincides with the usual expansion for the eigenvalue of the spheroidal harmonics [77]. Similarly, the asymptotic behavior for large $M\alpha$ can also be determined using monodromy techniques [87, 85].
On the other hand, the boundary conditions associated with the radial equation (1), i.e. purely ingoing wave at the event horizon and outgoing wave far away from it, satisfy the equation
$e^{\pi i\eta}=e^{\pi i\sigma}\frac{\sin\tfrac{\pi}{2}(\theta_{\star}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{\star}-\sigma)}\times\\ \frac{\sin\tfrac{\pi}{2}(\theta_{t}+\theta_{0}+\sigma)\sin\tfrac{\pi}{2}(\theta_{t}-\theta_{0}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{t}+\theta_{0}-\sigma)\sin\tfrac{\pi}{2}(\theta_{t}-\theta_{0}-\sigma)},$ (12)
@@
In this letter, we focus on the $\ell=m=1$ QNMs and explore the two-dimensional parameter space $(a/M,M\mu)$. We investigate the behavior of QNMs as the mass of the scalar field and the spin of the black hole change adiabatically away from the origin of the parameter space. At generic points, we follow the roots of the RH map with increments in $M\mu$ of order $10^{-3}$ and increments in $a/M$ of order $10^{-5}$. Near the exceptional point (see below) and near the extremal limit, we reduce the increments by 2 or 3 orders of magnitude.
We carry out the isomonodromic method using a Julia implementation of the Fredholm determinant formulation of $\tau_{V}$, truncated at $N_{f}=64$ levels. The determination of the parameter $\sigma$ employs a continued fraction approximant at $N_{c}=128$ levels. We use the ArbNumerics implementation for arbitrary precision arithmetic and dense matrix determinant calculations, working at $384$ digits. When possible, we used both small and large $t$ expansions to ensure numerical stability of at least $12$ digits.
@@
The fact that the $n=0$ modes become DMs for $M\mu>(M\mu)_{c}$ when the black hole is spinning sufficiently fast implies that they may not be the longest-living modes at every point of the parameter space. In fact, by following the $n=0$ and the $n=1$ QNMs throughout the parameter space, we have observed that their frequencies coincide at the critical point, which is thus understood as an exceptional point of a non-hermitian system. One of the exotic features of the exceptional point is shown in Fig. 3, where we plot both QNMs as a function of $M\mu$ for $a/M\lesssim(a/M)_{c}$ and $a/M\gtrsim(a/M)_{c}$. As $a/M$ increases through the critical value, we observe that the QNMs degenerate and transform into each other at $M\mu\simeq(M\mu)_{c}$.
The parameter space $(a/M,M\mu)$ is shown in Fig. 4, where the exceptional point is represented as a star. We have also identified a “curve of coexistence” emerging from the critical point, represented as a dashed curve, where the imaginary parts of the frequencies for the $n=0$ and the $n=1$ modes coincide. In the accompanying paper [86] we characterize the coexistence curve with greater detail. Note the instability curve [89], above which the longest-living QNMs are, in fact, unstable modes characterized by $\mathrm{Im}(\omega)>0$. When the QNMs are adiabatically deformed along a path that intersects the coexistence curve, such as the red path in Fig. 4, the longest-living mode switches to the second longest-living mode, and vice-versa.
@@
Geometric phase and hysteresis – So far we used $n=0$ and $n=1$ to refer to the spectrum of QNMs as ordered by their imaginary parts in the massless Schwarzschild configuration. Given the existence of an exceptional point, it may no longer be the case that following a mode adiabatically – for instance, $n=0$ or $n=1$, will yield the same results regardless of the path taken in parameter space. This dependence of the result of the deformation on the path raises the question of which mode will deform into the longest-living QNM in the near-extremal regime.
@@
To test this, we have adiabatically followed some of the modes around a closed path in parameter space containing the exceptional point. We have found that, even if the starting and ending points of the parameter space trajectory coincide, the final QNM frequency differs from the corresponding initial value, signalling the existence of a non-trivial geometric phase. Explicitly, consider the rectangle $ABCDA$ shown in Fig. 4, whose vertices $(a/M,M\mu)$ are $A=(0.999,0.45)$, $B=(0.999,0.3)$, $C=(0.9999,0.3)$ and $D=(0.9999,0.45)$. We start with the longest-living mode at $A$, whose frequency is $\omega_{\alpha}(A)=0.514974-0.026418i$. Changing the parameters continuously along the path $ABCD$, yields the longest-living mode at $D$, with frequency $\omega_{\alpha}(D)=0.500386-0.009456i$. On the other hand, deforming the QNM through the straight line $AD$ results in the second longest-living mode at $D$, with frequency $\omega_{\beta}(D)=0.514827-0.026255i$. If we repeat the procedure starting with the second longest-living mode at $A$, whose frequency is $\omega_{\beta}(A)=0.503306-0.032998i$, the results are reversed: the straight line leads to the longest-living mode at $D$, with frequency $\omega_{\alpha}(D)$, whereas the path that avoids the coexistence curve leads to the second longest-living mode at $D$, with frequency $\omega_{\beta}(D)$. To corroborate our findings, we also implemented Leaver’s method, as outlined in [90, 63, 91, 92], and obtained the same numerical results.
@@
Figure 3: The $n=0$ and $n=1$ QNMs as a function of $M\mu$ for $a/M\lesssim(a/M)_{c}$ (top panels) and $a/M\gtrsim(a/M)_{c}$ (bottom panels). At $M\mu\simeq(M\mu)_{c}$, the QNMs become degenerate and transform into each other, resembling the level crossing behavior observed in thermodynamic systems.
@@
Figure 4: Schematic representation of the parameter space associated with $\ell=m=1$ massive scalar perturbations in Kerr black holes. The star corresponds to the exceptional point where the $n=0$ and the $n=1$ QNMs are degenerate. The dashed line correponds to the coexistence curve where the decay rates of these QNMs are equal. Deformations of a QNM along the paths $AD$ (red) and $ABCD$ (dark green) yield different results, characterizing the ocurrence of hysteresis.
@@
We posit that the existence of the exceptional point and hysteresis phenomena we reported for Kerr black holes is analogous to the behavior of a thermodynamical system near a critical point. The critical point is the end of a curve of coexistence between phases, where there is a discontinuity of the free energy of the system. Likewise, in our coexistence curve the imaginary parts of the frequency match but the real part does not. In the thermodynamical system, continuing the curve beyond the critical point reveals a cross-over where the phases trasmute smoothly into one another, without any apparent discontinuity. We observe the same behavior by taking the $ABCDA$ path in Fig. 4, as if $\mathsf{K}(\omega)$ were a non-linear version of the transfer matrix of a statistical mechanical system.
Level-crossing and cross-overs are generic features of phase transitions in thermodynamics, as demonstrated in standard textbook treatments of the van der Waals fluid [94]. We have presented compelling evidence, through the specific example of massive scalar fields in Kerr black holes, that gravitational systems can display the same behavior.
Note added. After completing this work, we learned of a recent preprint by Motohashi [95], which interprets an avoided crossing near a degeneracy point as the mechanism underlying resonances between gravitational QNMs.

### v3

@@ Abstract
We employ the isomonodromic method to study linear scalar massive perturbations of Kerr black holes for generic scalar masses $M\mu$ and generic black hole spins $a/M$. We find that the longest-living quasinormal mode and the first overtone coincide for $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$. We also show that the longest-living mode and the first overtone change continuously into each other as we vary the parameters around the point of degeneracy, providing evidence for the existence of a geometric phase around an exceptional point. We interpret our findings through a thermodynamic analogy.
@@
$\partial_{u}\left[(1-u^{2})\partial_{u}S\right]+\left(\frac{m^{2}}{1-u^{2}}-c^{2}u^{2}-\Lambda\right)S=0,$ (3)
@@
$\begin{gathered}\theta_{0}=-m,\quad\theta_{t}=m,\quad\theta_{\star}=0,\quad\alpha=\sqrt{\omega^{2}-\mu^{2}},\\ t=-4a\alpha,\quad tc_{t}=\lambda+2(1-m)a\alpha+a^{2}\alpha^{2},\end{gathered}$ (5)
@@
$\begin{gathered}\theta_{0}=\frac{-i}{2\pi T_{-}}\left(\omega-m\Omega_{-}\right),\quad\theta_{t}=\frac{i}{2\pi T_{+}}\left(\omega-m\Omega_{+}\right),\\ \theta_{\star}=\frac{2iM(2\omega^{2}-\mu^{2})}{\alpha},\quad t=2i\alpha(r_{+}-r_{-}),\\ tc_{t}=\lambda+r_{+}^{2}\mu^{2}-(3a^{2}+r_{-}^{2}+3r_{+}^{2})\omega^{2}\\ \hskip 22.76228pt+i(r_{-}-r_{+}-2iam+2i(a^{2}+r_{+}^{2})\omega)\alpha\\ \hskip 42.67912pt+i\frac{M(2\omega^{2}-\mu^{2})}{\alpha}+\frac{M^{2}(2\omega^{2}-\mu^{2})^{2}}{\alpha^{2}}.\end{gathered}$ (6)
@@
The parameter $\sigma$ determined this way solves (7b). In fact, $\sigma$ parametrizes the complex phase of (8) as one continues it analytically over a path encompassing the singular points $z=0$ and $z=t$. In contrast, the interpretation of the monodromy parameter $\eta$ is quite involved [85]. For our purposes, we can just define the parameter by formally inverting the zero condition for $\tau_{V}$ (7a):
@@
where we have used $\theta_{0}$ and $\theta_{t}$ as defined in (5), and related the parameter $\sigma$ to the angular quantum number $\ell$. This equation relates the eigenvalues $\lambda$ and $\omega$ associated with the QNM problem. For $a\alpha\ll 1$, the expansion of $\lambda=\lambda(\omega)$ coincides with the usual expansion for the eigenvalue of the spheroidal harmonics [77]. Similarly, the asymptotic behavior for large $a\alpha$ can also be determined using monodromy techniques [87, 85].
Concerning the radial equation, QNMs are associated with boundary conditions which are purely ingoing at $r=r_{+}$ and purely outgoing at $r=\infty$. As discussed in more detail in [86], the mapping from $r$ in (1) to $z$ in (4) depends on $\alpha$, and for $\mathrm{Re}\,\alpha>0$ the QNM boundary condition translates to
$e^{\pi i\eta}=e^{-\pi i\sigma}\frac{\sin\tfrac{\pi}{2}(\theta_{\star}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{\star}-\sigma)}\times\\ \frac{\sin\tfrac{\pi}{2}(\theta_{t}+\theta_{0}+\sigma)\sin\tfrac{\pi}{2}(\theta_{t}-\theta_{0}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{t}+\theta_{0}-\sigma)\sin\tfrac{\pi}{2}(\theta_{t}-\theta_{0}-\sigma)},$ (12)
@@
In this letter, we investigate the $\ell=m=1$ QNMs across the two-dimensional parameter space $(a/M,M\mu)$, focusing on the question of what is the longest-living mode as the mass of the scalar field and the spin of the black hole vary. At generic points, we follow adiabatically the roots of the RH map with increments in $M\mu$ of order $10^{-3}$ and increments in $a/M$ of order $10^{-5}$. Near the exceptional point (see below) and near the extremal limit, we reduce the increments by 2 or 3 orders of magnitude.
We carry out the isomonodromic method using a Julia implementation of the Fredholm determinant formulation of $\tau_{V}$, truncated at $N_{f}=64$ levels. The determination of the parameter $\sigma$ employs a continued fraction approximant at $N_{c}=128$ levels. We use the ArbNumerics implementation for arbitrary precision arithmetic and dense matrix determinant calculations, working at $384$ digits. When possible, we used both small and large $t$ expansions to ensure numerical stability of at least $12$ digits. Numerical codes and datasets are publicly available [88].
@@
The fact that the $n=0$ modes become DMs for $M\mu>(M\mu)_{c}$ when the black hole is spinning sufficiently fast implies that they may not be the longest-living modes at every point of the parameter space. In fact, by following the $n=0$ and the $n=1$ QNMs throughout the parameter space, we have observed that their frequencies coincide at the critical point, which is thus understood as an exceptional point of a non-hermitian system. One of the exotic features of the exceptional point is shown in Fig. 3, where we plot both QNMs as a function of $M\mu$ for $a/M\lesssim(a/M)_{c}$ and $a/M\gtrsim(a/M)_{c}$. As the spin increases through the critical value, we observe that the QNMs degenerate and transform into each other. The parameter space $(a/M,M\mu)$ is shown in Fig. 4, where the exceptional point is represented as a star. We have also identified a “curve of coexistence” emerging from the critical point, represented as a dashed curve, where the imaginary parts of the frequencies for the $n=0$ and the $n=1$ modes coincide. In the accompanying paper [86] we characterize the coexistence curve with greater detail.
Note the curve separating the shaded (green) and white regions, defined by $\mathrm{Re}\,\alpha=0$ for the $n=0$ modes, above which the QNM boundary condition no longer follows Eq. (12). Extending the QNM analysis into the white region is beyond the scope of this letter, but some insights are discussed in [86]. When the QNMs are adiabatically deformed along a path that intersects the coexistence curve, such as the path $AD$ in Fig. 4, the longest-living mode switches to the second longest-living mode, and vice-versa.
@@
Geometric phase and hysteresis – So far we used $n=0$ and $n=1$ to refer to the spectrum of QNMs as ordered by their imaginary parts in the massless Schwarzschild configuration. The presence of an exceptional point means that following adiabatically either mode may yield different results depending on the path taken in parameter space. The path dependence establishes the existence of non-contractible curves in parameter space and ensures the existence of the exceptional point, whose numerical evidence we discussed above.
@@
To illustrate this, we have adiabatically followed some of the modes around a closed path in parameter space containing the exceptional point. We have found that, even if the starting and ending points of the parameter space trajectory coincide, the final QNM frequency differs from the corresponding initial value, signalling the existence of a non-trivial geometric phase. Explicitly, consider the rectangle $ABCDA$ shown in Fig. 4, whose vertices $(a/M,M\mu)$ are $A=(0.999,0.45)$, $B=(0.999,0.3)$, $C=(0.9999,0.3)$ and $D=(0.9999,0.45)$. We start with the longest-living mode at $A$, whose frequency is $\omega_{\alpha}(A)=0.514974-0.026418i$. Changing the parameters continuously along the path $ABCD$, and thus avoiding the coexistence curve, yields the longest-living mode at $D$, with frequency $\omega_{\alpha}(D)=0.500386-0.009456i$. On the other hand, deforming the QNM through the straight line $AD$, and thus crossing the coexistence curve, results in the second longest-living mode at $D$, with frequency $\omega_{\beta}(D)=0.514827-0.026255i$. If we repeat the procedure starting with the second longest-living mode at $A$, whose frequency is $\omega_{\beta}(A)=0.503306-0.032998i$, the results are reversed: the straight line leads to the longest-living mode at $D$, with frequency $\omega_{\alpha}(D)$, whereas the path that avoids the coexistence curve leads to the second longest-living mode at $D$, with frequency $\omega_{\beta}(D)$. To corroborate our findings, we also implemented Leaver’s method [88], as outlined in [90, 63, 91, 92, 93], and obtained the same numerical results.
@@
Figure 3: The $n=0$ and $n=1$ QNMs as a function of $M\mu$ for $a/M\lesssim(a/M)_{c}$ (left) and $a/M\gtrsim(a/M)_{c}$ (right). The arrows denote the direction of increasing $M\mu\in[0.0,0.8]$. At $\{a/M,M\mu\}=\{(a/M)_{c},(M\mu)_{c}\}$, the QNMs become degenerate and transform into each other, resembling the level crossing behavior of thermodynamical systems.
@@
Figure 4: Schematic representation of the parameter space associated with $\ell=m=1$ massive scalar perturbations in Kerr black holes. The star corresponds to the exceptional point where the $n=0$ and the $n=1$ QNMs are degenerate. The dashed line correponds to the coexistence curve where the decay rates of these QNMs are equal. Deformations of a QNM along the paths $AD$ and $ABCD$ yield different results, characterizing the ocurrence of hysteresis. The curve separating the shaded (green) and white regions is characterized by $\mathrm{Re}\,\alpha=0$ for the $n=0$ modes.
@@
We posit that the existence of the exceptional point and hysteresis phenomena we reported for Kerr black holes is analogous to the behavior of a thermodynamical system near a critical point. The critical point is the end of a curve of coexistence between phases, where there is a discontinuity of the free energy of the system. Likewise, in our coexistence curve the imaginary parts of the frequency match but the real part does not. In the thermodynamical system, continuing the curve beyond the critical point reveals a cross-over where the phases trasmute smoothly into one another, without any apparent discontinuity. We observe the same behavior by taking the $ABCD$ path in Fig. 4, as if $\mathsf{K}(\omega)$ were a non-linear version of the transfer matrix of a statistical mechanical system.
Level-crossing and cross-overs are generic features of phase transitions in thermodynamics, as demonstrated in standard textbook treatments of the van der Waals fluid [95]. We have presented compelling evidence, through the specific example of massive scalar fields in Kerr black holes, that gravitational systems can display the same behavior. Although localized in physical configuration space, these phenomena undoubtedly impact the ringdown phase and the overall behavior of massive fields when the relevant conditions are met. These topics present exciting opportunities for future research.
It is remarkable that the critical mass aligns with the regime where superradiant instabilities are most efficient, i.e. $(M\mu)_{c}\sim\mathcal{O}(1)$ [69]. In particular, putative bosonic condensates growing from these instabilities can emit gravitational waves through several channels (e.g. continuous emission arising from annihilation of the boson field or level transitions between states with different overtone numbers). Such processes (superradiant instabilities and gravitational wave emission) change the mass and the spin of the black hole, drawing paths in the parameter space (as in Fig. 4) that may cross a coexistence curve and encompass an exceptional point.
Note added. After completing this work, we learned of a recent preprint by Motohashi [96], which interprets an avoided crossing [97, 98, 99] near a degeneracy point as the mechanism underlying resonances between gravitational QNMs.